1-2 + 3 -4 + ....-98 + 99 equals
a) -50 b) -49 c) 0 d) 49 e) 50

Answer by Alan3354(6081)

1-2 + 3 -4 + ....-98 + 99 equals
a) -50 b) -49 c) 0 d) 49 e) 50
----------
There are 99 numbers
Pair them up:
1+99 = 100
-2-98 = -100
3+97 = 100
etc.
The pairs are alternately +100 and -100, until you get to -50. It has nothing to pair with, so the sum is
-50